Suppose that a company's equity is currently selling for $26.75 per share and that there are 4.1 million shares outstanding and 21 thousand bonds outstanding, which are selling at 98 percent of par. If the firm was considering an active change to their capital structure so that the firm would have a D/E of 1.6, which type of security (stocks or bonds) would they need to sell to accomplish this, and how much would they have to sell? (Round your intermediate ratio to 4 decimal places.)
Answer #1

Value of Equity = 26.75(4,100,000) = $109,675,000

Value of Debt = 980(21,000) = $20,580,000

Total Asset = 109,675,000 + 20,580,000 = $130,255,000

Required D/E = 1.6

So,

D/E = 1.6

Weight of Equity = 1/2.6 = 0.3846

So,

0.3846 = (109,675,000 - x)/(130,255,000 - x)

50,096,073 - 0.3846x = 109,675,000 - x

x = $96,813,336

So, firm needs to sell $96,813,336 of equity to maintain D/E of 1.6